# MedPromptExtract (Medical Data Extraction Tool): Anonymization and Hi-fidelity Automated data extraction using NLP and prompt engineering

### v1

@@ Abstract
A major roadblock in the seamless digitization of medical records remains the lack of interoperability of existing records. Extracting relevant medical information required for further treatment planning or even research is a time consuming labour intensive task involving the much valuable time of doctors. In this demo paper we present, MedPromptExtract an automated tool using a combination of semi supervised learning, large language models, natural language processing and prompt engineering to convert unstructured medical records to structured data which is amenable to further analysis.
1 Introduction
@@
Extracting relevant data from non digitized medical records can be a challenging task, moreover in low and middle income countries(LMICs) such data may not be available in interoperable formats Bajpai and Wadhwa (2021). Additionally, data may be required in retrievable formats for research purposesLehne et al. (2019). Discharge summaries, which provide comprehensive information about a patient’s history, treatment, and post-discharge care instructions, are valuable sources of such data Komenan et al. (2023). Anonymizing such records is essential to maintain confidentiality when used for research Rothstein (2010). Moreover, the implementation of artificial intelligence (AI) and machine learning, along with the adoption of paperless workflows, is crucial for transforming healthcare data and improving clinical decision-making.
@@
In this paper, we present MedPromptExtract a fully automated tool for effectively extracting data from discharge summaries while ensuring confidentiality.
2 Methodology
2.1 Dataset
@@
This study was a part of our long term goal to identify predictors of Acute Kidney Injury (AKI) in patients admitted to Kokilaben Dhirubhai Ambani Hospital (KDAH), Mumbai, India. We collected 914 discharge summaries retrospectively from patients likely to have AKI based on serum creatinine values, a biochemical product which is marker for AKI as per KDIGOKhwaja (2012) 2012 and AACCEl-Khoury et al. (2021) 2020 guidelines. All discharge summaries followed a standard format, facilitating automated data extraction. Ethical clearance was obtained from the Institutional Ethical Committee (IEC) of KDAH, Mumbai (IEC - A Code: 037/2021).
2.2 Anonymisation of discharge summaries
@@
We used EIGEN (Expert-Informed Joint Learning aGgrEatioN)Singh et al. (2023) for anonymization, leveraging large language models (LLMs) and semi-supervised learning techniques for high-fidelity information extraction from documents. We used Document Text Recognition (DocTR) Mindee (2021) to convert PDF files to JSON files for compatibility with EIGEN. EIGEN employs optical character recognition (OCR) and layout analysis to extract text and spatial structure from the discharge summaries. Labeling functions (LFs),which are are functions or heuristics that generate noisy or approximate labels for training data Ratner et al. (2016), were designed using positional heuristics to generate surrogate labels, significantly reducing annotation costs. Our training set consisted of 714 documents, of which only 20 were labeled, minimizing annotation costs.
Figure 1: Visualizing EIGEN
2.2.1 Model Implementation
We used LayoutLM Xu et al. (2020), a pre-trained deep neural network model, for information extraction from document images. The ”feature model” utilizes LayoutLM, while the ”graphical model” involves a probabilistic model that aggregates LF outputs to assign labels to unlabeled data. In our experiment, we had four classes to label: name, patient ID, location, and others. The batch size used was 32. The
@@ 2.2.1 Model Implementation
learning rate for the feature model and graphical model was $5\times 10^{-5}$ and 0.01 respectively. The optimizer used was AdamW. We used predicted class labels to anonymize discharge summaries by removing patient names and locations, replacing patient IDs with serial numbers. This process was conducted within the medical records department of KDAH, ensuring complete confidentiality. Figure 1 shows the visualization of bounding boxes for class labels after implementing EIGEN.
2.3 Extracting text from anonymised records
2.3.1 Preprocessing: Using Stopwords
@@
Figure 2: Designed prompt
@@
Figure 3: Flowchart of anonymisation and data extraction process
Feature Accuracy Sensitivity Specificity Precision
F1
Score
AUC
@@
Hyperkalemia Not computed as there were no true positives
Hypernatremia Not computed as there were no true positives
@@
Hypokalemia Not computed as there were no true positives
Hyponatremia Not computed as there were no true positives
@@
Table 1: Validation of model responses with clinicians responses
2.3.2 NLP Model Architecture
The extracted information was organized and stored in a structured DataFrame, with each stopword serving as a feature label for information following it.
2.4 Utilizing Prompt Engineering on Gemini to extract information
In this step, we analyzed free-flowing text under the heading ”course in hospital” to identify predictors of AKI. Nineteen features were identified through domain knowledge for extraction using the generative ability of LLMs and prompt engineering. A single prompt(Figure 2) was used to extract all clinical impressions, utilizing the Gemini APIPichai (2023). The API’s generative model, gemini-pro, was configured with temperature parameter set to 0 to ensure precise and deterministic outputs. Two iterations of model responses were subjected to inter-rater agreement using Kappa coefficient.
2.5 Validation with human annotations
Data extracted using LLM was validated against human annotations for 48 discharge summaries. Two clinicians annotated all summaries, and inter-rater agreement was assessed using the Kappa coefficient. Accuracy, precision, recall, F1 score and AUC were calculated using the ScikitLearn libraryPedregosa et al. (2011) in Python. Figure 3 illustrates the anonymisation and data extraction process of MedPromptExtract
3 Results
Out of the 914 discharge summaries processed, four were excluded initially due to format errors in the source PDFs received from the hospital. This anomaly was communicated to the hospital’s internal quality control cell. Upon running the prompt, it was observed that 58 summaries had an empty ”course in hospital” field, with two of them also included in the randomly selected annotations for human validation. Ultimately, outputs were obtained for 852 summaries, with 48 used for human validation. During annotation, it was found that seven of the 19 prompt questions had negative responses for all instances that is no true positives, leading to their exclusion from the final analysis.
Annotating a single summary took an average of 12 minutes, totaling 9.6 man-hours for 48 annotations. The compute time required for one iteration of 853 summaries was 100 minutes, equating to nearly 7 seconds per summary. Gemini’s API operates independently of the user’s local hardware, with computations conducted on their servers.
4 Limitations
@@
Our pipeline is custom-developed for the discharge summaries of the source hospital. While generalizable, it would require minor adjustments to the training of EIGEN and the prompts to be used for other similar usecases.
5 Conclusion
@@
Translating machine learning tools to real life hospital scenarios is a challenge given the practical constraints set by data quality and availability. This demo paper brings forth the ability of our model to perform automated anonymisation and data extraction with hi-fidelity utilizing very less labelled data, significantly reducing anonymisation costs, thus providing a good framework for digitization and further analysis of healthcare data. This method makes it possible to make data available in readily consumable format while maintaining anonymity. MedPromptExtract will be subsequently deployed in the source hospital and will be integrated with their EHR system for downstream applications.
6 Url for video

### v2

@@ Abstract.
A major roadblock in the seamless digitization of medical records remains the lack of interoperability of existing records. Extracting relevant medical information required for further treatment planning or even research is a time consuming labour intensive task involving expenditure of valuable time of doctors. In this demo paper we present, MedPromptExtract an automated tool using a combination of semi supervised learning, large language models, natural language processing and prompt engineering to convert unstructured medical records to structured data which is amenable for further analysis.
Digitizing Medical Records, Automated Anonymisation, Information Retrieval, Large Language Models, Prompt Engineering
conference: 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, Idaho, USAccs: Information systems Information extractionccs: Information systems Retrieval efficiency
1. Introduction
@@
Extracting relevant data from non digitized medical records can be a challenging task. Moreover, in low and middle income countries(LMICs) such data may not be available in interoperable formats (Bajpai and Wadhwa, 2021). Additionally, data may be required in retrievable formats for research purposes(Lehne et al., 2019). Discharge summaries (DS), which provide comprehensive information about a patient’s history, treatment, and post-discharge care instructions, are valuable sources of such data (Komenan et al., 2023). Anonymizing such records is essential to maintain confidentiality when used for downstream research applications (Rothstein, 2010). Moreover, the implementation of artificial intelligence (AI) and machine learning (ML), along with the adoption of paperless workflows, is crucial for transforming healthcare data and improving clinical decision-making. Our work is relevant to healthcare researchers working with unstructured data, especially in low resource settings, with the intention of using such data for tasks mentioned above.
@@
In this paper, we present MedPromptExtract a fully automated tool for effectively extracting data from DS while ensuring confidentiality.
2. Related Work
2.1. Automated Anonymisation
A comprehensive review by Meystre et al (Meystre et al., 2010) highlights common methods used to automate removal of protected health information (PHI) from medical text documents. These are observed to be either pattern recognition based or ML based. Pattern recognition uses rules, dictionaries and knowledge resources, while ML requires large sets of annotated records to effectively identify and remove such information. Work done by Deleger et al (Deleger et al., 2013) achieved de-identification of clinical records of 22 different formats with high precision and recall scores utilizing MITRE Identification Scrubber Toolkit (Aberdeen et al., 2010) which works on the procedure of tag-a-little, learn-a-little by annotating small sets of EHRs, training a model followed by hand correcting it which is labour intensive. More recent work such as (Mao and Liu, 2019) which utilizes BERT-CRF model for token level classification for anonymisation of medical records also requires fine tuning BERT on a large corpus.
2.2. Data Extraction from Structured Text
Most of the work for this purpose is focused on extracting data directly from Electronic Medical Records (EMRs) that is digitized sources of data. Recent work in this domain present similarities to our processes in the use of Natural Language Processes (NLP) such as use of regular expressions to extract data as reported in (C. Flint et al., 2021), whereas the older works mostly depended on use of Structured Query Language (SQL) as seen in (Kristianson et al., 2009). Data management through EMRs in LMICs such as India run on a hybrid system resulting in limited information available through EMRs while most of the patient details are entered in the patient’s offline records such as DS. Extraction of information related to medication from DS have been summarised in (Doan et al., 2012) and that for other clinical entities such as medical problems, tests, and treatments have also been explored in (Jiang et al., 2011). Both these works mention Name Entity Recognition (NER) using ensemble classifiers. This process requires annotated datasets which is a labour intensive process. In our work with, using advanced pdf processing tools and intelligently combining them with NLP tasks, we have achieved an efficient system requiring very little processing time and no prerequisites of large annotated data to extract relevant information.
2.3. Data extraction from unstructured text
The final step in our model includes the use of LLMs to extract custom information from free flowing text. This has also been attempted by many even prior to the era of LLMs, where NER has been extensively used. Most recently, work done by Adamson et al (Adamson et al., 2023) comes close to our objective to extract information in terms of binaries (yes/no, present/absent) or allocate the patient document to a particular category (malignant/non-malignant). In their process they first utilize NLP techniques to identify phrases of interest for extraction. The NLP derived features are inputs to a model which aims to assign the patient document to a particular class within a categorical variable. In their work they explored a number of deep learning architectures and achieved comparable results. Similar efforts were made by Hong et al (Hong et al., 2020) and Topaz et al (Topaz et al., 2016) where they used rule based system and pre-validated dictionary based NLP pipelines respectively to extract information from medical texts. However, they do acknowledge that such systems may fail where more complex information is considered such as description of medical events. In the same spectrum work by Lee et al (Lee et al., 2018) demonstrate that conditional random field (CRF) are more efficient for this purpose as far as semantics are concerned. In our work we demonstrate the utility of LLMs to extract context aware relevant information even without fine tuning on a corpus specific to this research problem or having to annotate the clinical text beforehand.
3. Methodology
3.1. Dataset
@@
This study was a part of our long term goal to identify predictors of Acute Kidney Injury (AKI) in patients admitted to Kokilaben Dhirubhai Ambani Hospital (KDAH), Mumbai, India. We collected 914 DS retrospectively from patients likely to have AKI based on serum creatinine values, a biochemical product which is marker for AKI as per KDIGO(Khwaja, 2012) 2012 and AACC(El-Khoury et al., 2021) 2020 guidelines. All DS followed a standard format, facilitating automated data extraction. Ethical clearance was obtained from the Institutional Ethical Committee (IEC) of KDAH, Mumbai (IEC - A Code: 037/2021).
3.2. Anonymisation of discharge summaries
@@
We used EIGEN (Singh et al., 2023) for anonymization, leveraging semi-supervised learning techniques for high-fidelity information extraction from documents. We used Document Text Recognition (DocTR) (Mindee, 2021) to convert PDF files to JSON files for compatibility with EIGEN. This model employs optical character recognition (OCR) and layout analysis to extract text and spatial structure from the DS. Labeling functions (LFs),which are are functions or heuristics that generate noisy or approximate labels for training data (Ratner et al., 2016), were designed using positional heuristics to generate surrogate labels, significantly reducing annotation costs. Our training set consisted of 714 documents, of which only 20 were labeled, minimizing annotation costs.
EIGEN uses LayoutLM (Xu et al., 2020), a pre-trained deep neural network, for information extraction from document images. Joint fine-tuning incorporates LFs into the training loop of LLMs, where the LLM known as the ”feature model” predicts class labels given words and their bounding boxes and probabilistic or ”graphical model” aggregates labels from LFs to assign labels to unlabeled data in the training set. In our experiment, we had four classes to label: name, patient ID, location, and others (which represents all the information we wish to extract). The batch size used was 32. The learning rate for the feature model and graphical model was $5\times 10^{-5}$ and 0.01 respectively. Data was trained on 4X48 GB GPU Server. The optimizer used was AdamW. We used predicted class labels to anonymize DS by removing patient names and locations, replacing patient IDs with serial numbers (Algorithm 1). This process was conducted within the medical records department of KDAH, ensuring complete confidentiality. Figure 1 shows the visualization of bounding boxes for class labels after implementing EIGEN.
Figure 1. Visualizing output of EIGEN
[…]
Visual representation of different class labels returned by EIGEN
@@
3.3. Extracting text from anonymised records
3.3.1. Preprocessing: Using Stopwords
@@
3.3.2. NLP Model Architecture
The extracted information was organized and stored in a structured DataFrame, with each stopword serving as a feature label for information following it (Algorithm 2).
@@ 3.3.2. NLP Model Architecture
3.4. Extracting unstructured text with Prompt engineering
In this step, we analyzed free-flowing text under the heading ”course in hospital” to identify predictors of AKI. Twelve features were identified through domain knowledge and interactions with nephrologists for extraction using the generative ability of LLMs and prompt engineering. A single prompt (Figure 2) was used to extract all clinical impressions, utilizing the Gemini API(Pichai, 2023). The API’s generative model, Gemini pro version 0.1.0, was configured with temperature parameter set to 0 to ensure precise and deterministic outputs. Two iterations of model responses were analyzed using Kappa coefficient for intra-model response agreement.
Figure 2. The designed prompt
[…]
The image list the 20 questions which were included in the prompt
3.5. Validation with human annotations
Two clinicians annotated 48 DS, and inter-rater agreement was assessed using the Kappa coefficient. Data extracted using LLM was validated against this as ground truth by calculating accuracy, precision, recall, F1 score and AUC using the ScikitLearn library(Pedregosa et al., 2011) in Python. Figure 3 illustrates the pipeline of MedPromptExtract.
Figure 3. Workflow of MedPromptExtract Tool
[…]
A flowchart with three columns detailing the three main pipelines used to anonymize DS and extract data from them, also representing number of samples processed at each stage
4. Results
Anonymisation pipeline took 46 minutes to run that is approximately 3 seconds for one DS. Out of the 914 DS processed, four were excluded initially due to format errors in the source PDFs received from the hospital. Our NLP pipeline took 2.5 minutes to run, resulting in 0.2 seconds per DS. Upon running the prompt using Gemini pro, it was observed that 58 summaries had an empty ”course in hospital” field, with two of them also included in the randomly selected annotations for human validation. Ultimately, outputs were obtained for 852 summaries, with 48 used for human validation. The compute time required for 852 summaries was 100 minutes, equating to nearly 7 seconds per summary. On the other hand time taken for Human Annotations was 9.6 man-hours for 48 annotations at the rate of 12 minutes per document.
Table 1. Validation of model responses with clinicians’ responses
Feature Acc. Sen. Spec. Pre. F1 AUC
@@ 4. Results
Note: Acc. = Accuracy, Sen. = Sensitivity, Spec. = Specificity,
Pre. = Precision, F1 = F1 Score, AUC = Area Under the Curve
5. Discussion and Limitations
In our approach we demonstrate high fidelity data extraction in response to custom prompts without any fine tuning of the model used or annotations done on the input text. Seven out of twelve features in our tested prompt achieve area under the curve of more than 0.9, the remaining 5 did not perform so well; the reasons for this are discussed here.
To investigate this, randomly selected DS giving positive and negative responses to these five questions were checked by experts to qualitatively assess its responses. It was noted that the model was giving contextual responses to certain questions. For example, even if the DS did not mention AKI, but discussed a rise in creatinine values then the responses awarded as presence of AKI was positive by the model. Similarly in questions like whether general anaesthesia was given or whether patient was admitted to the ICU, responses by the model were contextually derived that is patient having undergone any kind of surgery was assumed to have undergone general anaesthesia, this however is not always true. Other questions which suffered from such assumption was Oxygen Saturation being equated with any kind of breathing issues. The question on consultation by nephrologist was a unique case as model was unable to match human annotations as in most summaries the name of the nephrologist was referred to rather than the speciality, resulting in the doctor annotations identifying it correctly (as they were from the same hospital), but model not being able to match it.
@@ 5. Discussion and Limitations
A limitation is that our pipeline is custom-developed for the DS of the source hospital. While generalizable, it would require minor adjustments to the training of EIGEN and the prompts to be used for other similar use-cases.
6. User Interface
The user interface (UI) of MedPromptExtract has been developed for this tool and is demonstrated in the video attached. The interface is simple and clear with intuitive navigation options and options for saving/downloading outputs at every point as per the need. The UI also provides the option to edit the prompt to derive information that may be different from what we have demonstrated here. For the purpose of demonstration, in the video, we show both the options of using our predefined prompt or using a different prompt and visualizing the output of both. We demonstrate processing of a single DS for the purpose of understanding and thereafter we also display the ability of our UI to process bulk summaries with processing 20 at the same time which took less than 5 minutes to process.
7. Conclusion
@@
Translating ML tools to real life hospital scenarios is a challenge given the practical constraints set by data quality and availability. This demo paper brings forth the ability of our methodology to perform automated anonymisation and data extraction with hi-fidelity utilizing very less labelled data, significantly reducing anonymisation costs, thus providing a good framework for digitization and further analysis of healthcare data. MedPromptExtractwill be subsequently deployed in the source hospital and will be integrated with their EHR system for downstream applications.
8. Url for video